Find equivalent mass in reation C2O4-2 -->CO2Please give the explanation on how to find n factor and equivalent mass in these type of reations

Grade:12

1 Answers

Vikas TU
7 years ago
Eq. mass is given as => Molecular mass/Valence factor
or moles x Valence Fcator.
Hence to calculate the valence factor first find the oxidation state for the atom in both sides that is in reactant and Product side.
In Reacn.  C2O4-2                      →                 CO2
x = +6/2 = +3                                       x = +4
Valnce factor for Crabon = (4-3)*2 = 2.